OpenStudy (anonymous):

how do you find y"(x) for y = the square root of 5x^2 + 4 the 5x^2 + 4 is all under the square root

OpenStudy (amistre64):

chain rule ...

OpenStudy (amistre64):

\[\frac d{dx}[f(x)]^n=n[f(x)]^{n-1}~*~\frac d{dx}f(x)\]

OpenStudy (amistre64):

look like the second round will involve a product rule

OpenStudy (anonymous):

But how do you get rid of the square root in order to do the first derivative

OpenStudy (amistre64):

it helps to recall your algebra :)\[\sqrt{M}=M^{1/2}\]

OpenStudy (anonymous):

so, is it (5x^2 + 4) ^1/2 ?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

\[(5x^2+4)^{1/2}\to 5x*(5x^2+4)^{-1/2}\]

OpenStudy (anonymous):

therefore, it would be: 1/2 (5x^2 + 4 ) ^-1/2 ?

OpenStudy (amistre64):

dont forget to pop out that chain rule of 5x^2+4

OpenStudy (anonymous):

where are you getting the first 5x from that is outside the parentheses ? I am confused on that

OpenStudy (anonymous):

Becuase I thought for the first derivative it is just the derivative of: (5x^2 + 4) ^1/2 ... --> 1/2(5x^2 +4) ^-1/2?

OpenStudy (amistre64):

the chain rule: let u = 5x^2 + 4; u' = 2*5x \[u^{1/2}\to \frac{u'}{2u^(1/2)}\to \frac{\cancel2*5x}{\cancel2\sqrt{5x^2+4}}\]

OpenStudy (amistre64):

whenever you have a function wrapped inside of another function, you have to apply the chain rule ... which is hard to grasp at first since they dont tell you that you always pop out a chain rule even when you were learning the process

OpenStudy (anonymous):

oh okay I gotcha, so, can we back this up for a moment... In order to find the first derivative of this problem, y = the square root of 5x^2 + 4.... you put 1/2(5x^2 + 4) ^-1/2 ? and that is the first derivative?

OpenStudy (amistre64):

thats incomplete ... so no

OpenStudy (anonymous):

So is it 1/2(10x)^-1/2? or am I doing that derivative incorrectly

OpenStudy (amistre64):

do it in steps .. you seem to be confusing a few tings at the moment

OpenStudy (amistre64):

\[\frac d{dx}[f(x)]^n=n[f(x)]^{n-1}~*~\frac d{dx}f(x)\]

OpenStudy (amistre64):

let f(x) = 5x^2 + 4 the derivative of f(x) is therefore 10x and n=1/2 fill in the parts \[\frac d{dx}[f(x)]^n=n[f(x)]^{n-1}~*~\frac d{dx}f(x)\] \[\frac d{dx}[5x^2+4]^{1/2}=\frac12[5x^2+4]^{-1/2}~*~10x\]

OpenStudy (anonymous):

So, do we need to simplify that for the first derivative to be complete?

OpenStudy (amistre64):

you can if you want .... 10/2 = 5 is what i did up top

OpenStudy (amistre64):

\[5x~(5x^2+4)^{-1/2}\] now this is a product rule ... and the chain rule will still apply for this one too \[5'x~(5x^2+4)^{-1/2}~+~5x~(5x^2+4)'^{-1/2}\] \[5~(5x^2+4)^{-1/2}~-~5x~\frac12(5x^2+4)'^{-3/2}*10x\] \[5~(5x^2+4)^{-1/2}~-~25x(5x^2+4)^{-3/2}\] factor out the -3/2 stuff \[(5~(5x^2+4)^{4/2}~-~25x)(5x^2+4)^{-3/2}\] \[\frac{5~(5x^2+4)^{4/2}~-~25x}{(5x^2+4)^{3/2}}\] and the rest is simple algebra stuff

OpenStudy (amistre64):

mighta factored wrong up top :) make that a 1/2 exponent instead

OpenStudy (amistre64):

\[\frac{5~(5x^2+4)^{1/2}~-~25x}{(5x^2+4)^{3/2}}\] thats feels better